December 17, 2014 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, December 17, 2014
11:00 am, EDT
JLab: CEBAF Center, F326

Connection Using Bluejeans

  1. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 630804895.
  2. To join via a Web Browser, go to the page [1]
  3. To join via phone, use one of the following numbers and the Conference ID: 630804895
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
  5. More information on connecting to bluejeans is available.


  1. Announcements
  2. Status Update
  3. Calibration status/updates
    1. Calorimetry
    2. Tracking
    3. Timing
      1. TOF Matching (Paul)
        TOF Matching (Paul)
  4. Calibration Skims
  5. Simulations
  6. Data Monitoring
    1. Quick summary by Kei
    2. Commissioning Software
    3. Run Browser, Plot Browser, and Time Series webpages
  7. Peak hunting / higher-level analyses
  8. AOB


Attendees: Mark D., Mark I., Dave M., Paul M., Will M., Curtis M., Kei M., Zisis P., Lubomir P., Eric P., Andrei S., Matt S., Elton S., ST, Justin S., Beni Z.

  • Status of ccdb:
    • Entries for rollover/tframe for TDC has been added to ccdb. The relevant parameters will also be in the data stream (at the beginning of the run? how do we handle the multiple files per run?)
    • Placeholder for v_eff for TOF scintillators has been added.
  • BCAL
    • SiPM gains based on Andrei's initial analysis
      • No measured constants for ~50% of the channels in the BCAL due to too low gain, threshold effects -- put in rough guess
      • Will M. has reconstructed π0 peak at ~160 MeV
      • Mark D. added scale factors to move peak down to nominal mass.
    • Next step is to add attenuation lengths to the database (the table already exists)
      • Measurements from beam data give values ~520 cm, whereas earlier studies indicated values closer to 380 cm using phototube readout
    • Mark D. has seen some strange 2-peak behavior in some TDC channels in the BCAL -- these double peaks do not necessarily show up in the same TDC from one run to the next.
    • Elton has been working on a draft document describing the BCAL calibration constants: [2]
  • FCAL
    • Matt S. is working on calibration using π0s; he's working on filtering out the noise by looking at 3 photon events with the aim to cut on an ω(782) peak
  • Tracking
    • Lubomir has been analyzing a large (~6 hour) data set with B=0 and FCAL+TOF in the trigger for FDC alignment. He was able to repeat results obtained for Package 3 in the EEL.
    • Vlad B. has performed a second pass of the strip-to-strip gain calibration.
    • ST has been working on the CDC alignment using a long cosmic run (8.7 M triggers!)
  • Skims: the 2-track skim has been very useful so far. There was some discussion about whether other skims such as a 4 track skim (for searching for detached vertices) or a π0 skim would be useful.
    • Mark D. suggested a skim for studying pedestals.
  • Paul M. showed plots of track matching to the time-of-flight. These show large variations from counter to counter that are not understood at the moment.
  • Simulation: new simulations are needed for the current conditions (for example the active target was moved downstream). It would be nice to model the tilt of the start counter as well. ST will work on the xml geometry.
  • Kei presented a status report about the offline monitoring
    • Of 2760 files processed, 37 had bad alloc errors, 137 had "insufficient buffer size" errors (these are actually caused by corrupted events), 76 had "unknown module type" errors, and 55 had some other JANA crash.
    • The question was asked whether or not we should continue with the weekly jobs -- feeling was yes. Also brought up was the idea of a weekly analysis test train (maybe looking at the first one or two files in each run).
    • Kei noticed that there was a discrepancy in file sizes between the tape silo and the gluon raid for 17 files out of 5000.
  • Higher level analysis will be done using rest files at a later date (after an initial round of calibrations).